zepdrix (zepdrix):

Consider the forced Van der Pol oscillator equation: \[\huge\bf\sf \ddot x-\mu(1-x^2)\dot x+x\quad=\quad A \sin(\omega t)\] 1. Express the equation as a system of differential equations.

zepdrix (zepdrix):

WUT? D: HOW?

zepdrix (zepdrix):

I only see one independent variable. Am I trying to write it as a sytem of `x` and `t` maybe? Hmmm this is weird..

OpenStudy (abb0t):

@shamil98

OpenStudy (shamil98):

@abb0t I don't know this stuff. Why don't you show us your brilliance? hm?

zepdrix (zepdrix):

lol XD

zepdrix (zepdrix):

I don't need to solve it or anything :d I'm using approximation methods and computers for that. Just trying to figure out what the "system" should look like.. hmm

OpenStudy (anonymous):

Set \(y_1=x\) and \(y_2=x'\), so that you see that \(y_1'=y_2\) and \(y_2'=x''\). So then your original equation can be written as \[y_2'-\mu(1-y_1^2)+y_1=A\sin\omega t\] and you have the system \[\begin{cases}y_1'=y_2\\y_2'=\mu(1-y_1)^2-y_1+A\sin\omega t\end{cases}\]

zepdrix (zepdrix):

Ooo nice :D thanks